### v4

@@ Abstract
In this work, we developed a new Bayesian method for variable selection in function-on-scalar regression (FOSR). Our method uses a hierarchical Bayesian structure and latent variables to enable an adaptive covariate selection process for FOSR. Extensive simulation studies show the proposed method’s main properties, such as its accuracy in estimating the coefficients and high capacity to select variables correctly. Furthermore, we conducted a substantial comparative analysis with the main competing methods, the BGLSS (Bayesian Group Lasso with Spike and Slab prior) method, the group LASSO (Least Absolute Shrinkage and Selection Operator), the group MCP (Minimax Concave Penalty), and the group SCAD (Smoothly Clipped Absolute Deviation). Our results demonstrate that the proposed methodology is superior in correctly selecting covariates compared with the existing competing methods while maintaining a satisfactory level of goodness of fit. In contrast, the competing methods could not balance selection accuracy with goodness of fit. We also considered a COVID-19 dataset and some socioeconomic data from Brazil as an application and obtained satisfactory results. In short, the proposed Bayesian variable selection model is highly competitive, showing significant predictive and selective quality.
keywords:
Bayesian inference, functional data, functional data analysis, variable selection, latent variable, function-on-scalar regression
@@ 1 Introduction
Linear regression models are already well consolidated, and well understood in the statistical literature [22, 34, 38], with most models developed considering cross-sectional rather than functional data. When the data are functions, the standard linear regression models, generally, cannot be used directly because they require building of a design matrix that contains not only the covariates information but also other important elements, such as the basis functions coming from the expansions of the functional data. This strategy can be observed in some works such as in [37], [7] and [3]. Functional regression [32, Chapters 12 to 17] then emerged to circumvent this problem, being one of the areas of functional data analysis that has received a lot of attention, both in applied studies and in methodological developments [32, 16, 26, 7, 33, 31].
It is worth to mention three types of functional regression modeling as described in [32]: i) the functional predictor regression (scalar-on-function), in which covariates and coefficients are functional but the response is not; ii) the regression with functional response (function-on-scalar), where the covariates are vectors and the coefficients are functional; iii) and the regression that models the association through functional coefficients, between functional covariates and a functional response (function-on-function).
Bearing in mind the context of functional regression and the growing demand for selection models in this area, herein we focus on a regression problem of the function-on-scalar type and propose a Bayesian selection model of covariates. In what follows, we describe in more details the proposed function-on-scalar regression (FOSR) model (ii). For a description of the other types of modeling (i and iii described above), see [32].
In function-on-scalar regression, we consider a set of scalar predictor variables $x_{li}$, where $i$ is the index linked to the $i$th curve and $l$ is the one associated with the $l$th covariate. Then, considering a linear structure, we have the following model:
@@
So, it is also necessary to impose some kind of regularization to avoid overfitting. Notice that the intercept and variable coefficients are functional. The component $\epsilon_{i}(t)$ represents the error curve, or residual curve. It is often assumed that the error curves are independent, and come from the same Gaussian process with a mean of zero and a given covariance structure.
In many cases, in addition to the goal on prediction, the aim is to establish a good estimation of the functional coefficients in (1), assessing whether $\beta_{l}(t)=0$ for all $t$ for some $l$, or even checking whether a given functional coefficient is null only at specific points, meaning if $\beta_{l}(t)=0$ only for some values of $t$.
@@
The literature on variable selection in functional regression analysis is still not dominant. However, there is strong evidence that this area has grown rapidly in recent years, especially in the last decade. Some examples of variable selection in functional regression can be found in [25], which studied the group SCAD regularization for the selection of functional regressors, [27] and [15], which generalized LASSO to the case of scalar covariates and functional response, among others [10, 1, 13, 24, 8]. A related but different problem in functional data analysis is the selection of basis functions for functional data representation (smoothing). In this context, one can use penalty-based techniques such as LASSO [36] and Bayesian LASSO [29] or latent variable approaches as in [21], [2] and [35].
There is, specifically, a growing interest in research problems on FOSR. Some works from the last decade explore the potential of FOSR under a Bayesian perspective, as can be seen in [28] and in [19], as well as [20], which additionally handles variable selection. However, the frequentist approach is predominant when it comes to variable selection in FOSR. Several studies with a focus on variable selection in FOSR were developed and applied to genetic datasets [37, 11, 3, 30], although there are other areas of applications, such as for Alzheimer’s data [6], among others [7, 20].
@@
This work proposes a new Bayesian approach to select covariates in a FOSR model by associating a Bernoulli latent variable to each functional coefficient. This makes it possible to assign zero to specific model functional coefficients with a positive probability. From a hierarchical structure, considering a linear combination of basis functions to represent each functional coefficient, we obtain the full conditional distributions to build a Gibbs sampler to sample from the posterior distribution (R code implementation is available upon request). Unlike other Bayesian approaches that commonly use non-informative priors on the last layer, we consider an exponential prior to the coefficient variance component to prevent problems involving colinearity among variables.
@@
This paper is organized as follows. Section 2 presents the proposed Bayesian model for variable selection in FOSR, while Section 3 shows the design and results of several numerical experiments involving the proposed methodology. Also, in Section 3, we present a comparative study between the proposed model and the methods group LASSO, group SCAD, group MCP and BGLSS. In Section 4, we conduct a study to evaluate the performance of the proposed model in a functional regression problem involving some socioeconomic data and COVID-19 data from the Federal District and Brazilian states. Finally, Section 5 provides some general conclusions about this work.
2 Proposed Bayesian model for variable selection in functional regression
[…]
@@ 2 Proposed Bayesian model for variable selection in functional regression
where $Z_{l}$ is a Bernoulli latent variable that will select or not the $l$th covariate. While $y_{ij}$ represents the value of the $i$th curve at point $t_{ij}$ (meaning $y_{ij}=y_{i}(t_{ij})$), the partial functional coefficients associated with the covariates are represented by $\beta_{l}(.)$’s, just as $\beta_{0}(.)$ represents the intercept. Finally, $\epsilon_{ij}=\epsilon_{i}(t_{ij})$ is the random component that represents the error measurements with a normal distribution with zero mean and variance equal to $\sigma^{2}$.
For a given fixed curve $i$, the vector with the final functional coefficients of this functional linear regression is given by $\boldsymbol{\nu(.)}=(\nu_{1}(.),\;\nu_{2}(.),\dots,\;\nu_{p}(.))^{{}^{\prime}}=(Z_{1}\beta_{1}(.),\;Z_{2}\beta_{2}(.),\dots,$ $\;Z_{p}\beta_{p}(.))^{{}^{\prime}}$. For the sake of brevity and to avoid confusion in the notation, the components $\nu_{l}(.)$’s will be called functional coefficients, while the $\beta_{l}(.)$’s will be called partial functional coefficients.
@@
where the $b_{kl}$’s represent the coefficients of the basis expansion for the $l$th partial functional coefficient, and the $b_{k0}$’s are the coefficients of the basis expansion applied on the functional intercept.
Hence, developing the Bayesian model for adaptive variable selection through a hierarchical structure is possible. First, however, some definitions should be established for ease of understanding. Let $\boldsymbol{\theta}=(\theta_{1},\;\theta_{2},\dots,\;\theta_{p})^{{}^{\prime}}$ and $\boldsymbol{Z}=(Z_{1},\;Z_{2},\dots,\;Z_{p})^{{}^{\prime}}$. While $\boldsymbol{\theta}$ and $\boldsymbol{Z}$ are vectors of dimension $(p\times 1)$, there are two other vectors of higher dimension $(Kp\times 1)$ that are: $\boldsymbol{\tau^{2}}=(\boldsymbol{\tau^{2}_{.1}}^{{}^{\prime}},\;\boldsymbol{\tau^{2}_{.2}}^{{}^{\prime}},\dots,\;\boldsymbol{\tau^{2}_{.p}}^{{}^{\prime}})^{{}^{\prime}}$, where $\boldsymbol{\tau^{2}_{.l}}=(\tau^{2}_{1l},\tau^{2}_{2l},\dots,\tau^{2}_{Kl})^{{}^{\prime}}$ and $\boldsymbol{b}=(\boldsymbol{b_{.1}}^{{}^{\prime}},\;\boldsymbol{b_{.2}}^{{}^{\prime}},\dots,\;\boldsymbol{b_{.p}}^{{}^{\prime}})^{{}^{\prime}}$, with $\boldsymbol{b_{.l}}=(b_{1l},b_{2l},\dots,b_{Kl})^{{}^{\prime}}$. Note that there is also a vector that contains all $b_{k0}$’s, say, $\boldsymbol{b_{0}}=(b_{10},\;b_{20},\dots,\;b_{K0})^{{}^{\prime}}$. It is worth mentioning that the intercept will be estimated separately; therefore, a prior will not be assigned to the $b_{k0}$’s. Finally, consider $\boldsymbol{y}=(\boldsymbol{y_{1.}}^{{}^{\prime}},\;\boldsymbol{y_{2.}}^{{}^{\prime}},\dots,\;\boldsymbol{y_{m.}}^{{}^{\prime}})^{{}^{\prime}}$, where $\boldsymbol{y_{i.}}=(y_{i1},y_{i2},\dots,y_{in_{i}})^{{}^{\prime}}$. Then, we have:
@@
A sample from the posterior distribution of $\boldsymbol{Z}$, $\boldsymbol{b}$ and $\sigma^{2}$ is obtained via the Gibbs sampler described in Section 2.2.

### v5

@@ Abstract
In this work, we developed a new Bayesian method for variable selection in function-on-scalar regression (FOSR). Our method uses a hierarchical Bayesian structure and latent variables to enable an adaptive covariate selection process for FOSR. Extensive simulation studies show the proposed method’s main properties, such as its accuracy in estimating the coefficients and high capacity to select variables correctly. Furthermore, we conducted a substantial comparative analysis with the main competing methods: the BGLSS (Bayesian Group Lasso with Spike and Slab prior) method, the group LASSO (Least Absolute Shrinkage and Selection Operator), the group MCP (Minimax Concave Penalty), and the group SCAD (Smoothly Clipped Absolute Deviation). Our results demonstrate that the proposed methodology is superior in correctly selecting covariates compared with the existing competing methods while maintaining an excellent level of goodness of fit. In contrast, the competing methods could not balance selection accuracy with goodness of fit. We also considered a COVID-19 dataset and some socioeconomic data from Brazil as an application and obtained sound results. In short, the proposed Bayesian variable selection model is highly competitive, showing significant predictive and selective quality.
Keywords: Bayesian inference, functional data, functional data analysis, variable selection, latent variable, function-on-scalar regression
@@ 1 Introduction
Linear regression models are well-established and thoroughly studied in the statistical literature (Kutner et al., 2004; Rencher and Schaalje, 2008; Weisberg, 2013), but most have been developed for cross-sectional data rather than functional data. When the data are functions, standard linear regression models are generally unsuitable, as they require constructing a design matrix that incorporates not only covariate information but also additional elements, such as basis functions derived from functional data expansions. This approach has been explored in studies such as Wang et al. (2007), Chen et al. (2016), and Barber et al. (2017). To address these limitations, functional regression models (Ramsay and Silverman, 2005, Chapters 12 to 17) were introduced, becoming a key area of functional data analysis that has attracted considerable attention in both applied and methodological research (Ramsay and Silverman, 2005; Horváth and Kokoszka, 2012; Meyer et al., 2015; Chen et al., 2016; Reiss et al., 2017; Qi and Luo, 2018).
Among the various functional regression models described in Ramsay and Silverman (2005), this paper focuses on function-on-scalar regression (FOSR), where the response is functional, covariates are scalar, and coefficients are functional. We propose a Bayesian model for parameter inference and covariate selection within the FOSR framework. Under a linear structure, the FOSR model is expressed as:
@@
where, $y_{i}(t)$ is the value of the $i$th response functional evaluated at point $t$, $x_{li}$ is the value of the $l$th scalar predictor corresponding to the $i$th curve. Notice that the intercept and variable coefficients are functional. The component $\epsilon_{i}(t)$ represents the error curve. It is often assumed that the error curves are independent and come from the same Gaussian process with a mean of zero and a given covariance structure. In many cases, in addition to the goal of prediction, the aim is to estimate the functional coefficients in (1) while imposing some regularization to avoid overfitting. This can be done by assessing whether $\beta_{l}(t)=0$ only for some values of $t$ or whether $\beta_{l}(t)=0$ for all $t$. Here, we focus on the latter to perform variable selection.
@@
The literature on variable selection in functional regression analysis is still not dominant. However, there is strong evidence that this area has grown rapidly in recent years, especially in the last decade. Some examples of variable selection in functional regression can be found in Matsui and Konishi (2011), which studied the group SCAD regularization for the selection of functional regressors, Mingotti et al. (2013) and Hong and Lian (2011), which generalized LASSO to the case of scalar covariates and functional response, among others (Fan and Li, 2004; Aneiros et al., 2011; Gertheiss et al., 2013; Ma et al., 2013; Collazos et al., 2016). A related but different problem in functional data analysis is the selection of basis functions for functional data representation (smoothing). In this context, one can use penalty-based techniques such as LASSO (Tibshirani, 1996) and Bayesian LASSO (Park and Casella, 2008) or latent variable approaches as in Kuo and Mallick (1998), Anselmo et al. (2005) and Sousa et al. (2023).
There is growing interest in research problems focusing on FOSR. Some studies over the last decade have explored its potential from a Bayesian perspective (Montagna et al., 2012; Kowal, 2021; Kowal and Bourgeois, 2020), with the latter also addressing variable selection. However, the frequentist approach remains predominant for variable selection in FOSR, with applications spanning genetic studies (Wang et al., 2007; Fan and Reimherr, 2017; Barber et al., 2017; Parodi and Reimherr, 2018), Alzheimer’s research (CAI et al., 2022), and other domains (Chen et al., 2016).
@@
This work proposes a new Bayesian approach to select covariates in an FOSR model by associating a Bernoulli latent variable with each functional coefficient. This makes it possible to assign zero to specific model functional coefficients with positive probability. From a hierarchical structure, considering a linear combination of basis functions to represent each functional coefficient, we obtain the full conditional distributions to build a Gibbs sampler to sample from the posterior distribution (R code implementation is available upon request). Unlike other Bayesian approaches that commonly use non-informative priors on the last layer, we use an exponential prior on the coefficient variance component to prevent problems of collinearity among variables. We also conduct an extensive comparative analysis between our proposed methodology and each of the following competing methods: group LASSO (Yuan and Lin, 2006), group SCAD (Wang et al., 2007), group MCP (Huang et al., 2012) and BGLSS (Xu and Ghosh, 2015).
@@
This paper is organized as follows. Section 2 presents the proposed Bayesian model for variable selection in FOSR, while Section 3 shows the design and results of several numerical experiments involving the proposed methodology. Also, in Section 3, we present a comparative study between the proposed model and the competing methods. In Section 4, we conduct a study to evaluate the performance of the proposed model in a functional regression problem involving some socioeconomic data and COVID-19 data from the Federal District and Brazilian states. Finally, Section 5 provides some general conclusions about this work.
2 Proposed Bayesian model for variable selection in FOSR
For the introduction of the proposed Bayesian model, let us consider $p$ covariates, $m$ curves with $n_{i}$ observations at the points $t_{ij}\in A\subseteq\mathbb{R}$, in which $i\in\{1,\;2,\dots,\;m\}$ and $j\in\{1,\;2,\dots,\;n_{i}\}$. So, we assume that:
@@ 2 Proposed Bayesian model for variable selection in FOSR
where $Z_{l}$ is a Bernoulli latent variable that selects or not the $l$th covariate. While $y_{ij}$ represents the value of the $i$th curve at point $t_{ij}$ (meaning $y_{ij}=y_{i}(t_{ij})$), the partial functional coefficients associated with the covariates are represented by $\beta_{l}(.)$’s, just as $\beta_{0}(.)$ represents the intercept. Finally, $\epsilon_{ij}=\epsilon_{i}(t_{ij})$ is the random component that represents the error measurements with a normal distribution with zero mean and variance equal to $\sigma^{2}$.
For a given fixed curve $i$, the vector with the final functional coefficients of this functional linear regression is given by $\boldsymbol{\nu(.)}=(\nu_{1}(.),\;\nu_{2}(.),\dots,\;\nu_{p}(.))^{{}^{\prime}}=(Z_{1}\beta_{1}(.),\;Z_{2}\beta_{2}(.),\dots,$ $Z_{p}\beta_{p}(.))^{{}^{\prime}}$. For the sake of brevity and to avoid confusion in the notation, the components $\nu_{l}(.)$’s are called functional coefficients, while the $\beta_{l}(.)$’s are called partial functional coefficients.
@@
where the $b_{kl}$’s represent the coefficients of the basis expansion for the $l$th partial functional coefficient, and the $b_{k0}$’s are the coefficients of the basis expansion applied to the functional intercept.
Hence, it is possible to develop a Bayesian model for adaptive variable selection with a hierarchical structure. First, however, some definitions should be established to facilitate understanding. Let $\boldsymbol{\theta}=(\theta_{1},\;\theta_{2},\dots,\;\theta_{p})^{{}^{\prime}}$ and $\boldsymbol{Z}=(Z_{1},\;Z_{2},\dots,\;Z_{p})^{{}^{\prime}}$. While $\boldsymbol{\theta}$ and $\boldsymbol{Z}$ are vectors of dimension $(p\times 1)$, there are two other vectors of higher dimension $(Kp\times 1)$ that are: $\boldsymbol{\tau^{2}}=(\boldsymbol{\tau^{2}_{.1}}^{{}^{\prime}},\;\boldsymbol{\tau^{2}_{.2}}^{{}^{\prime}},\dots,\;\boldsymbol{\tau^{2}_{.p}}^{{}^{\prime}})^{{}^{\prime}}$, where $\boldsymbol{\tau^{2}_{.l}}=(\tau^{2}_{1l},\tau^{2}_{2l},\dots,\tau^{2}_{Kl})^{{}^{\prime}}$ and $\boldsymbol{b}=(\boldsymbol{b_{.1}}^{{}^{\prime}},\;\boldsymbol{b_{.2}}^{{}^{\prime}},\dots,\;\boldsymbol{b_{.p}}^{{}^{\prime}})^{{}^{\prime}}$, with $\boldsymbol{b_{.l}}=(b_{1l},b_{2l},\dots,b_{Kl})^{{}^{\prime}}$. Note that there is also a vector that contains all $b_{k0}$’s, say, $\boldsymbol{b_{0}}=(b_{10},\;b_{20},\dots,\;b_{K0})^{{}^{\prime}}$. It is worth mentioning that the intercept is estimated separately; therefore, a prior is not assigned to the $b_{k0}$’s. Finally, consider $\boldsymbol{y}=(\boldsymbol{y_{1.}}^{{}^{\prime}},\;\boldsymbol{y_{2.}}^{{}^{\prime}},\dots,\;\boldsymbol{y_{m.}}^{{}^{\prime}})^{{}^{\prime}}$, where $\boldsymbol{y_{i.}}=(y_{i1},y_{i2},\dots,y_{in_{i}})^{{}^{\prime}}$. Then, we have:
@@
A sample from the posterior distribution of $\boldsymbol{Z}$, $\boldsymbol{b}$ and $\sigma^{2}$ is obtained via the Gibbs sampler described in Subsection 2.1.